# Combinative Matching for Geometric Shape Assembly

### v1

@@ 2 Related Work
Complementary matching for assembly. A recent work by Lu et al[24] presents the concept of a primal-dual descriptor to reflect viewpoint-dependent characteristics for surface matching. Similarly to our motivation, they intend to capture the essence of complementary geometry arising from fracture assembly. However, focusing on the characteristics of a local surface from two different directions, inward and outward, they separate a descriptor into primal and dual ones, and train them to align by intercrossing in matching, i.e., encouraging the primal descriptor of one part to resemble the dual descriptor of the other part in matching. Despite a similar motivation, the primal-dual matching method implements the geometric complementarity of mating parts simply by switching two distinct descriptors in matching, and train them to resemble in the primal-dual pair between mating parts; there is no clear distinction between the primal and dual descriptors in terms of their roles and effects. This is clearly different from our approach that distinguishes the descriptor for surface shape, which is to be identical between mating parts, from that for volumetric occupancy, which is to be opposite between mating parts. As will be discussed in our experimental section and supplementary, the primal-dual descriptors fail to capture the interlocking properties of mating parts, and our combinative matching clearly outperforms in matching performance.
@@ 3.1 Combinative Matching
$\mathcal{L}_{\text{s}}\propto\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\log\!\Bigl{(}\sum_{j\in\mathcal{E}_{\text{p}}(i)}e^{(d_{ij}^{\text{p}}-\Delta_{\text{p}})^{2}}\cdot\sum_{k\in\mathcal{E}_{\text{n}}(i)}e^{(\Delta_{\text{n}}-d_{ik}^{\text{n}})^{2}}\Bigr{)}\right],$ (2)
where $d_{ij}^{\text{p}}=\|\hat{\mathbf{F}}_{\text{s},i}^{\text{P}}-\hat{\mathbf{F}}_{\text{s},j}^{\text{Q}}\|_{2}$ represents the L2 distances between shape features in the embedding space, where $\hat{\mathbf{F}}$ denotes the L2-normalized features, and $\mathcal{E}_{\text{p}}(i),\mathcal{E}_{\text{n}}(i)$ are positive/negative correspondences for index $i$, and $\Delta_{\text{p}},\Delta_{\text{n}}$ are margin hyperparameters. This formulation encourages distance for positive matches to be close to the threshold $\Delta_{\text{p}}$ while pushing the distance for negative ones to exceed $\Delta_{\text{n}}$, similar to the conventional approaches that identify identical surface geometries.
@@
$\mathcal{L}_{\text{o}}\propto\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\log\!\Bigl{(}\sum_{j\in\mathcal{E}_{\text{p}}(i)}e^{(s_{ij}^{\text{p}}-\Delta_{\text{p}})^{2}}\cdot\sum_{k\in\mathcal{E}_{\text{n}}(i)}e^{(\Delta_{\text{n}}-s_{ik}^{\text{n}})^{2}}\Bigr{)}\right],$ (3)
where $s_{ij}^{\text{p}}=\|\hat{\mathbf{F}}_{\text{o},i}^{\text{P}}+\hat{\mathbf{F}}_{\text{o},j}^{\text{Q}}\|_{2}\approx\cos(\mathbf{F}_{\text{o},i}^{\text{P}},\mathbf{F}_{\text{o},j}^{\text{Q}})$ represents the cosine similarity measures in the occupancy embedding space. It is worth noting that while conventional circle loss typically uses distance metrics to bring positive pairs closer together, our approach leverages cosine similarity to explicitly encourage opposite occupancy between positive pairs. This approach treats occupied-unoccupied pairs as positive matches, encouraging their descriptors to be opposite, while penalizing non-matching pairs. Thus, the model learns to identify complementary volumes that interlock, rather than matching identical geometries.
@@ 4.1 Dataset and Evaluation Metrics
Evaluation Metrics. We use the evaluation metrics used in PMTR [16] to validate our method: (1) RMSE between ground truth and predicted rotation and translation parameters, (2) CoRrespondence Distance (CRD), the average distance between positive matches on mating surfaces, and (3) Chamfer Distance (CD) between the input and model-predicted assemblies. Note that CRD provides a more reliable measure than RMSE(R,T) as CRD directly assesses assembly quality while RMSE(R,T) measures relative pose differences without explicitly considering alignment accuracy. Following Lee et al[16], our evaluation is performed with relative poses between parts instead of absolute ones to solely focus on the assembly, not absolute positioning.
@@ 4.5 Comparison with State of the Arts
We further evaluate our method on multi-part assembly with additional metrics, Part Accuracy (PA) [19, 16], and compare the results in Tab. 4 and Fig. 8, where ours consistently shows superior numbers compared to baselines. The results confirm that, unlike methods that rely solely on visual cues [34, 16], our combinative matching enables more reliable shape assembly, as evident from Figs. 7 and 8. We refer to the supplementary for additional qualitative results.
@@ B.5 Details on training objectives
Circle loss [40]. In the definition of circle loss provided in the main paper, we omit hyperparameter $\gamma$ used in the original circle loss formulation, for brevity in our demonstration. The definition of $\mathcal{L}_{\text{circle}}$ is formulated as
$\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\log\left(\sum_{j\in\mathcal{E}_{\text{p}}(i)}e^{\gamma(\phi^{i,j}-\Delta_{\text{p}})^{2}}\cdot\sum_{k\in\mathcal{E}_{\text{n}}(i)}e^{\gamma(\Delta_{\text{n}}-\phi^{i,k})^{2}}\right)\right],$ (10)
where $\gamma$ scales the sharpness of the exponential terms, amplifying or reducing the emphasis on outliers and enabling more stable training and $\phi^{i,j}$ computes similarity or dissimilarity between a pair of given features.
Point matching loss [35]. We now detail the point matching loss $\mathcal{L}_{\text{p}}$ which is jointly used alongside the combinative matching objectives. The point matching loss is a negative log-likelihood loss on predicted match probabilities $\mathbf{Z}\in\mathbb{R}^{(N+1)\times(M+1)}$, the dual-normalized assignment matrix with dustbin, the output from the Optimal Transport layer [35]. Given a set of ground-truth correspondences $\mathcal{C}$, and the sets of unmatched points, $(\mathcal{I}^{\text{P}})^{\prime}=\{x:x\in[N]\wedge x\notin\mathcal{I}^{\text{P}}\}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ (\mathcal{I}^{\text{Q}})^{\prime}=\{x:x\in[M]\wedge x\notin\mathcal{I}^{\text{Q}}\}$ where $K]=\{1,\dots,K\}$, the point matching loss is defined as

### v2

@@ 2 Related Work
Complementary matching for assembly. A recent work by Lu et al. [24] presents the concept of a primal-dual descriptor to reflect viewpoint-dependent characteristics for surface matching. Similarly to our motivation, they intend to capture the essence of complementary geometry arising from fracture assembly. However, focusing on the characteristics of a local surface from two different directions, inward and outward, they separate a descriptor into primal and dual ones, and train them to align by intercrossing in matching, i.e., encouraging the primal descriptor of one part to resemble the dual descriptor of the other part in matching. Despite a similar motivation, the primal-dual matching method implements the geometric complementarity of mating parts simply by switching two distinct descriptors in matching, and train them to resemble in the primal-dual pair between mating parts; there is no clear distinction between the primal and dual descriptors in terms of their roles and effects. This is clearly different from our approach that distinguishes the descriptor for surface shape, which is to be identical between mating parts, from that for volumetric occupancy, which is to be opposite between mating parts. As will be discussed in our experimental section and supplementary, the primal-dual descriptors fail to capture the interlocking properties of mating parts, and our combinative matching clearly outperforms in matching performance.
@@ 3.1 Combinative Matching
$\mathcal{L}_{\text{s}}=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\log\!\Bigl(\sum_{j\in\mathcal{E}_{\text{p}}(i)}e^{\alpha_{ij}(d_{ij}^{\text{p}}-\Delta_{\text{p}})}\cdot\sum_{k\in\mathcal{E}_{\text{n}}(i)}e^{\beta_{ik}(\Delta_{\text{n}}-d_{ik}^{\text{n}})}\Bigr)\right],$ (2)
where $d_{ij}^{\text{p}}=\|\hat{\mathbf{F}}_{\text{s},i}^{\text{P}}-\hat{\mathbf{F}}_{\text{s},j}^{\text{Q}}\|_{2}$ represents the L2 distances between shape features in the embedding space, where $\hat{\mathbf{F}}$ denotes the L2-normalized features, and $\mathcal{E}_{\text{p}}(i),\mathcal{E}_{\text{n}}(i)$ are positive/negative correspondences for index $i$, and $\Delta_{\text{p}},\Delta_{\text{n}}$ are margin hyperparameters. The positive and negative weights are computed as $\alpha_{ij}=\gamma[d_{ij}^{\text{p}}-\Delta_{\text{p}}]_{+}$ and $\beta_{ik}=\gamma[\Delta_{\text{n}}-d_{ik}^{\text{n}}]_{+}$, with scale factor $\gamma$. This formulation encourages the distance for positive matches to fall below the threshold $\Delta_{\text{p}}$ while pushing the distance for negative ones to exceed $\Delta_{\text{n}}$, similar to conventional approaches that identify identical surface geometries.
@@
$\mathcal{L}_{\text{o}}=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\log\!\Bigl(\sum_{j\in\mathcal{E}_{\text{p}}(i)}e^{\alpha_{ij}(s_{ij}^{\text{p}}-\Delta_{\text{p}})}\cdot\sum_{k\in\mathcal{E}_{\text{n}}(i)}e^{\beta_{ik}(\Delta_{\text{n}}-s_{ik}^{\text{n}})}\Bigr)\right],$ (3)
where $s_{ij}^{\text{p}}=\|\hat{\mathbf{F}}_{\text{o},i}^{\text{P}}+\hat{\mathbf{F}}_{\text{o},j}^{\text{Q}}\|_{2}\approx\cos(\mathbf{F}_{\text{o},i}^{\text{P}},\mathbf{F}_{\text{o},j}^{\text{Q}})$ represents the cosine similarity measures in the occupancy embedding space, and the positive and negative weights are computed as $\alpha_{ij}=\gamma[s_{ij}^{\text{p}}-\Delta_{\text{p}}]_{+}$ and $\beta_{ik}=\gamma[\Delta_{\text{n}}-s_{ik}^{\text{n}}]_{+}$. It is worth noting that while conventional circle loss typically uses distance metrics to bring positive pairs closer together, our approach leverages cosine similarity to explicitly encourage opposite occupancy between positive pairs. This approach treats occupied-unoccupied pairs as positive matches, encouraging their descriptors to be opposite, while penalizing non-matching pairs. Thus, the model learns to identify complementary volumes that interlock, rather than matching identical geometries.
@@ 4.1 Dataset and Evaluation Metrics
Evaluation Metrics. We use the evaluation metrics used in PMTR [16] to validate our method: (1) RMSE between ground truth and predicted rotation and translation parameters, (2) CoRrespondence Distance (CRD), the average distance between positive matches on mating surfaces, and (3) Chamfer Distance (CD) between the input and model-predicted assemblies. Note that CRD provides a more reliable measure than RMSE(R,T) as CRD directly assesses assembly quality while RMSE(R,T) measures relative pose differences without explicitly considering alignment accuracy. Following Lee et al. [16], our evaluation is performed with relative poses between parts instead of absolute ones to solely focus on the assembly, not absolute positioning.
@@ 4.5 Comparison with State of the Arts
We further evaluate our method on multi-part assembly [^4] As noted in Sec. 4.1, our comparisons are conducted on the volume-constrained Breaking Bad dataset. The evaluation results on the vanilla Breaking Bad dataset are provided in our supplementary material. with additional metrics, Part Accuracy (PA) [19, 16], and compare the results in Tab. 4 and Fig. 8, where ours consistently shows superior numbers compared to baselines. The results confirm that, unlike methods that rely solely on visual cues [34, 16], our combinative matching enables more reliable shape assembly, as evident from Figs. 7 and 8. We refer to the supplementary for additional qualitative results.
@@ B.5 Details on training objectives
Circle loss [40]. The definition of $\mathcal{L}_{\text{circle}}$ is formulated as
$\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\log\left(\sum_{j\in\mathcal{E}_{\text{p}}(i)}e^{\alpha_{ij}(\phi^{i,j}-\Delta_{\text{p}})}\cdot\sum_{k\in\mathcal{E}_{\text{n}}(i)}e^{\beta_{ik}(\Delta_{\text{n}}-\phi^{i,k})}\right)\right],$ (10)
where the positive and negative weights are defined as $\alpha_{ij}=\gamma[\phi^{i,j}-\Delta_{\text{p}}]_{+}$ and $\beta_{ik}=\gamma[\Delta_{\text{n}}-\phi^{i,k}]_{+}$, where $\gamma$ controls the sharpness of the exponential terms, amplifying or reducing the emphasis on outliers and enabling more stable training. And $\phi^{i,j}$ denotes the similarity (or dissimilarity) between a given feature pair. In our implementation, the circle loss uses margin parameters of $\Delta_{\text{p}}=0.1$ and $\Delta_{\text{n}}=1.4$, with $\gamma=24$.
Point matching loss [35]. We now detail the point matching loss $\mathcal{L}_{\text{p}}$ which is jointly used alongside the combinative matching objectives. The point matching loss is a negative log-likelihood loss on predicted match probabilities $\mathbf{Z}\in\mathbb{R}^{(N+1)\times(M+1)}$, the dual-normalized assignment matrix with dustbin, the output from the Optimal Transport layer [35]. Given a set of ground-truth correspondences $\mathcal{C}$, and the sets of unmatched points, $(\mathcal{I}^{\text{P}})^{\prime}=\{x:x\in[N]\wedge x\notin\mathcal{I}^{\text{P}}\}\penalty 10000\ \text{and}\penalty 10000\ (\mathcal{I}^{\text{Q}})^{\prime}=\{x:x\in[M]\wedge x\notin\mathcal{I}^{\text{Q}}\}$ where $K]=\{1,\dots,K\}$, the point matching loss is defined as